I need help with addition and subtraction of fractions. Example 8 7/8 - 3 4/9
I need to show my work, can you help

Answer by Alan3354(69443)

I need help with addition and subtraction of fractions. Example 8 7/8 - 3 4/9
8 7/8 - 3 4/9
Change to improper fractions, not mixed
71/8 - 31/9
Find a common denominator, 8*9 = 72
Change to common denominators
639/72 - 248/72
= (639 - 248)/72
= 391/72
---------
You can change it to a mixed fraction, but there's no reason to do that.